OpenStudy (anonymous):

Find the general solution \[{dP \over dt} = 0.08P\left(1-{P \over 1000} \right) -15\]

OpenStudy (anonymous):

I started by multpling and factoring the right side

OpenStudy (anonymous):

then it became seperable and I ended up with \[\int{25 \over P - 250} - {25 \over P - 750} dP = - \int dt\] after partial fractions

OpenStudy (anonymous):

\[25 \ln {|P - 250| \over |P - 750|} = -t + c\] how do I solve this thing for P?

OpenStudy (anonymous):

Do you need to make it into a P= something form? Because that one up there is already a considerable final answer, in my opinion.

OpenStudy (anonymous):

It's a population growth equation, i need to use it to solve for some populations